OpenStudy (anonymous):

Which statement best describes the function: f(x)=x^3-x^2-9+9 ??? a. It is a many-to-one function. b. It is not a function. c. It is a one-to-one function. d. It fails the vertical line test. My answer: C. It is a one-to-one function Am i right?? Please explain!! Thanks! :)

OpenStudy (anonymous):

Hi! thanks for coming!!! Am i right? :)

OpenStudy (dumbcow):

yes,a cubic function is one-to-one

OpenStudy (anonymous):

So my answer is correct since it is a cubic function?

OpenStudy (anonymous):

and is it a cubic function because it is -9 +9??

OpenStudy (anonymous):

Or is there a different explanation of a cubic function?

OpenStudy (dumbcow):

well -9+9 = 0 ?

OpenStudy (anonymous):

is that what a cubic function is?

OpenStudy (experimentx):

x^3-x^2-9+9 = x^3 - x^2 = x^2(1-x) let it be zero, you will get roots ... x=0 or x=1 f(0) = 0 f(1) = 0

OpenStudy (anonymous):

it pass vertical line test. but horizontal line test fails so not one to one function